Please Help!!!!
x-y=1
x^2+y^2=5

The point (x y) is in the 3rd quadrant of the xy- plane and satisfies the equations above. What is value of y?

Answer by stanbon(75887)

x-y=1
x^2+y^2=5
----
x = y+1
-----
Substitute for "x" and solve for "y"::
(y+1)^2 + y^2 = 5
2y^2 + 2y + 1 = 5
2y^2 + 2y - 4 = 0
----------------------
y^2 + y - 2 = 0
-----
Factor:
(y+2)(y-1) = 0
----
y = 1 or y = -2
--------------------
Solve for "x"::
x = y+1
If y = 1, x = 2
If y = -2, x = -1
---------------
Cheers,
Stan H.
---------------